# UniGen: A Unified Framework for Textual Dataset Generation via Large Language Models

### v3

@@ Abstract
Large Language Models (LLMs) such as GPT-4 and Llama3 have significantly impacted various fields by enabling high-quality synthetic data generation and reducing dependence on expensive human-generated datasets. Despite this, challenges remain in the areas of generalization, controllability, diversity, and truthfulness within the existing generative frameworks. To address these challenges, this paper presents UniGen, a comprehensive LLM-powered framework designed to produce diverse, accurate, and highly controllable datasets. UniGen is adaptable, supporting all types of text datasets and enhancing the generative process through innovative mechanisms. To augment data diversity, UniGen incorporates an attribute-guided generation module and a group checking feature. For accuracy, it employs a code-based mathematical assessment for label verification alongside a retrieval-augmented generation technique for factual validation. The framework also allows for user-specified constraints, enabling customization of the data generation process to suit particular requirements. Extensive experiments demonstrate the superior quality of data generated by UniGen, and each module within UniGen plays a critical role in this enhancement. Additionally, UniGen is applied in two practical scenarios: benchmarking LLMs and data augmentation. The results indicate that UniGen effectively supports dynamic and evolving benchmarking, and that data augmentation improves LLM capabilities in various domains, including agent-oriented abilities and reasoning skills.
@@ 1 Introduction
Large Language Models (LLMs) such as GPT-4 [1], Claude [2], and Llama3 [3] have demonstrated excellent performance across various professional domains, including medical [4, 5], educational [6], software engineering [7], and social sciences [8], as well as in LLM-based agent applications [9, 10, 11]. Given their superior generative capabilities, researchers are naturally inclined to explore effective methods for utilizing these models in synthetic data generation [12, 13, 14]. The primary objective is to produce high-quality, cost-effective datasets, thereby reducing the reliance on expensive human labor. Furthermore, data generated by LLMs can be employed for data augmentation [15], dynamic evaluation [12, 13], enhancing model self-alignment [16], and more.
Despite the advancements in LLM-generated data [12, 13, 14, 17, 18], which have significantly improved the data generation pipeline and reduced the human cost, some challenges remain, particularly in terms of (1) Generalization and Controllability: Most of existing frameworks directly modify data items in original datasets in specific ways based on fixed principles [13, 14] (e.g., add additional context or shuffle the order of the options), which may constrain the generalization of the generated data as they do not modify the nature of the data items like the scenarios within items. Moreover, many of them are also limited to particular dataset formats or types [15, 12], such as multiple-choice or mathematically-oriented datasets (e.g., GSM8K [19]). Additionally, the lack of provisions for incorporating external constraints, like specific user requirements (e.g., users may specify the length of generated text), restricts their controllability during generation. (2) Diversity and Truthfulness: Prior efforts always overlook the need to ensure some quality aspects of the datasets like diversity and truthfulness. For instance, the direct application of LLMs for dataset generation often leads to replication and low diversity, as LLMs may output the same answers when faced with semantically similar input. Furthermore, the propensity of LLMs to produce hallucinations [20, 21] can introduce factual inaccuracies, potentially degrading model performance when such datasets are used for training or fine-tuning.
Figure 1: Our proposed UniGen for dataset generation via LLMs.
@@
DyVal [12] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
DyVal 2 [13] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
S3Eval [22] $\checkmark$ $\checkmark$ $\checkmark$
Yu et al. [15] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Chung et al. [23] $\checkmark$ $\checkmark$
Fan et al. [24] $\checkmark$ $\checkmark$ $\checkmark$
Jandaghi et al. [25] $\checkmark$ $\checkmark$
Wang et al. [14] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
[…]
To address these challenges, this paper puts forward UniGen (as shown in Figure 1), a unified and LLM-powered framework designed to generate a dataset. UniGen ensures the generalization, diversity, truthfulness, and controllability simultaneously of the generation process, compared to previous studies (as shown in Table 1). UniGen accepts all kinds of text datasets and generates high-quality datasets based on various modules. To enrich the diversity of the generated datasets, UniGen employs a range of strategies, including various hyperparameter settings, attribute-guided generation, and group checking. To guarantee the truthfulness of the generated datasets, we propose a code-based mathematical assessment to detect and rectify potentially incorrect labels. Additionally, we adopt a Retrieval-Augmented Generation (RAG)-based validation method to check the factuality of generated statements to ensure their truthfulness. UniGen integrates constraints input to align with user specifications to enhance user control over the dataset generation process. Furthermore, by employing attribute-guided generation and difficulty enhancement, we enable the generation of data covering a wide range of topics while providing users with controllable difficulty levels.
@@
    We introduce a unified framework, UniGen, specifically designed for generating textual datasets via LLMs. UniGen accepts the original dataset, dataset description, and user constraints, as well as integrates different modules to ensure diversity, truthfulness, and controllability during generation.
  • $\triangleright$
@@
    We carry out extensive experiments to assess the effectiveness of UniGen, covering aspects such as data characterization, module efficacy, human evaluation, error analysis, and cost analysis. The results affirm that UniGen is proficient in dataset generation and suggests promising directions for future research.
  • $\triangleright$
@@
    Furthermore, we delve into two potential applications of UniGen —benchmarking LLMs and data augmentation. Our findings provide several key insights. For example, I) Most LLMs struggle with math-oriented datasets generated by UniGen (e.g., GSM8K). II) The benchmark performance of LLMs varies significantly across datasets generated by different LLMs. III) LLMs’ capabilities across various aspects (e.g., agent-related abilities, reasoning skills) can be improved by fine-tuning based on the generated data. IV) A potential improvement of data augmentation still exists in knowledge-intensive datasets.
  • $\triangleright$
    Based on the observations and findings presented, Appendix A discusses the limitations of the current framework for dataset generation and proposes potential improvement measures for future studies. These enhancements are considered from multiple perspectives, including error analysis, downstream applications, and LLM alignment.
2 Related Work
Benchmarking and Evaluating LLMs.
Owing to the remarkable capabilities of LLMs, benchmarking these models is essential for a deeper understanding of both general and specialized domains [26]. The evaluation of LLMs encompasses a wide range of fields, initiating with core NLP tasks such as sentiment analysis [27, 28], text classification [29, 30], and natural language inference [31]. A holistic evaluation framework, the HELM benchmark, has been proposed by Liang et al. [32], laying the groundwork for comprehensive assessments. Additionally, the application of LLMs spans diverse sectors [33], including computational social science [34], legal analytics [35, 36, 37], and psychological studies [38, 8]. Furthermore, several benchmarks have been designed to scrutinize trustworthiness dimensions such as safety and privacy in LLMs [21, 39, 40, 41, 42, 43]. Static benchmarks are susceptible to data contamination, wherein developers might incorporate benchmark datasets into the training data to artificially enhance performance. To mitigate this, flexible protocols for dynamic evaluation have been advanced, exemplified by the recent initiatives DyVal [12] and DyVal 2 [13]. Additionally, Fan et al. [24] introduced NPHardEval, featuring monthly updated datasets. The S3Eval framework, a scalable evaluation suite for LLMs, was conceptualized by [22]. Moreover, some benchmarks adopt methodologies where LLMs function as evaluators (e.g., LLM-as-a-judge) [44, 45, 46], with AlignBench proposing a multi-dimensional assessment using this approach [44].
Synthetic Data by LLMs.
LLMs have demonstrated an impressive capacity for data generation, leading to their application in creating synthetic datasets for pretraining and fineu’runing, replacing the labor-intensive processes of manual data scraping and selection [47]. Distinct from earlier methods that focus on traditional language models [48], LLMs offer enhanced prospects for producing high-quality synthetic data across a wide spectrum of applications, such as multilingual QA [49], chatbot conversation [50] and data diversity augmentation [51, 23].
The concept of synthetic benchmarks takes a step further by demanding that the LLM-generated data be diverse accurate and systematically challenging. For instance, Wang et al. [14] devised a framework that enhances the evolution of benchmarks by applying six reframing techniques on existing datasets. Wei et al. [52] employed GPT-4 to create LongFact, comprising extensive QA pairs that serve as a benchmark for evaluating long-form factual content. Moreover, synthetic benchmarks have also been constructed in evaluating LLM emergent capabilities such as trustworthiness [21], tool usage [9, 53] and persona-based conversation [25]. Our research advances synthetic benchmark generation by developing a paradigm that integrates multiple plug-and-play modules into LLM dataset creation, leveraging emergent capabilities by various prompting methods (e.g., self-evaluation [54]) to produce data items with high-quality. Recently, in response to concerns about the quality of synthetic datasets, Dekoninck et al. [17] conducted comprehensive experiments to evaluate the diversity and fidelity of synthetic data produced by LLMs, while Dekoninck et al. [18] introduced a new inference framework, model arithmetic, to control the content generated by LLMs.
3 UniGen Framework
Figure 2: The architecture of UniGen.
@@ 3 UniGen Framework
In this section, we will introduce the proposed UniGen, a unified framework for dataset generation. UniGen consists of four modules (as shown in Figure 2) including framework input, generation hint, internal evaluation, and post-processing. Formally, consider an original dataset $\mathcal{D}$, the proposed framework $\mathcal{F}$, which operates by iteratively sampling subsets $\mathcal{S}_{i}$ from $\mathcal{D}$ (i.e., example selection for few-shot learning in subsection 3.2). For each subset, $\mathcal{F}$ applies transformations based on the dataset’s description $\mathcal{M}(\mathcal{D})$ and a set of constraints $\mathcal{C}$. The final generated dataset, $\mathcal{D}_{\text{gen}}$, is accumulated over $N$ iterations: $\mathcal{D}_{\text{gen}}=\bigcup_{i=1}^{N}\mathcal{F}(\mathcal{S}_{i},\mathcal{M}(\mathcal{D}),\mathcal{C})$. During generation, the objectives of UniGen focus on maximizing the generalization, controllability, diversity, and truthfulness of the generated dataset.
To ensure optimal functionality, UniGen integrates four main modules:
  • $\triangleright$
    Framework Input: This module captures the basic information about the target dataset (e.g., data examples, dataset description) and any explicit constraints specified by users for data generation.
  • $\triangleright$
    Generation Hint: Designed to steer the generation process of LLMs, this module ensures that the generated dataset closely mirrors the original dataset and aligns with user specifications.
  • $\triangleright$
    Internal Evaluation: This module evaluates and refines the generated dataset. It plays a crucial role in enhancing dataset quality, including assessing the truthfulness of the generated data items.
  • $\triangleright$
    Post-Processing: Used to perform additional operations on the generated dataset, this module adds flexibility to the framework, adapting it to various application needs.
@@ 3.1 Framework Input
The input for UniGen comprises three components: base dataset, dataset description, and generation constraints: The base dataset is provided in a standardized JSON format, which may include text with a label or standalone text (e.g., “text with a label” or “single text”). The dataset description articulates the specifics of the base dataset at a high level, furnishing foundational guidance for the LLM to synthesize a dataset analogous to the original. While optional, the generation constraints [55] specify fine-grained conditions under which the LLM operates during dataset generation. For instance, constraints might stipulate that “Do not generate text longer than 20 words” or “Include an emoji in each generated sample”, thereby restricting specific conditions of the synthetic dataset.
@@ 3.2 Generation Hint
Few-Shot Learning. The base dataset typically comprises hundreds of data items; however, incorporating all these items directly into the prompt may result in an excessively long context that could obscure the comprehension capabilities of LLMs and incur substantial costs [56]. To mitigate these challenges, few-shot learning techniques are employed for dataset generation [57, 58]. Within UniGen, two principal methods are utilized to select few-shot learning examples. The first method involves a random sampling from the base dataset, effectively reducing both generation time and associated costs. The second method focuses on enhancing the diversity of examples, thereby guiding LLMs to generate as varied a dataset as possible. Specifically, UniGen initially encodes all data items using OpenAI’s text-embedding-ada-002 [59] to create an embedding list. Subsequently, a clustering algorithm (e.g., K-means [60]) is applied to form $n$ clusters, where $n$ represents the desired number of examples. One example is randomly selected from each cluster, yielding a set of $n$ diverse examples.
Diversity Setting. To augment the diversity of the generated data, we implement two strategies: (1) Hyperparameter Setting. The content generated by LLMs is influenced by various factors, with hyperparameters such as temperature, top-k, and top-p being crucial. To maximize the diversity of the dataset, we manipulate these hyperparameters, particularly the temperature settings. (2) Attribute-Guided Generation. Drawing on insights from prior research [9, 15], we formalize the attribute-guided text generation process for LLMs. Let $\mathcal{A}=\{a_{1},a_{2},\dots,a_{n}\}$ be a set of attributes, such as "economics" and "sports", intended to guide the generation process. We model the generation process as a function where the output text $y$ is a function of the input prompt $x$ and a vector of attributes $\mathbf{a}\in\mathcal{A}$. The generation process can be expressed as $y=P(x,\mathbf{a})$, where $P$ represents the generation model of the LLM, and $x$ is the input prompt. To implement this, we employ two distinct strategies: the first involves directly incorporating user-input customized attributes, and the second requires asking LLMs to extract necessary attributes from given data examples (the prompt template is shown in Appendix H). (3) Group Checking. To ensure diversity among the generated items, a similarity matrix is employed to identify and filter out pairs of data items exhibiting high similarity. Further details on this process are provided in subsection 3.4.
@@ 3.3 Internal Evaluation
Overall Quality Assessment and Enhancement. After obtaining the raw generated data, it’s important to enhance their overall quality as during the generation, LLMs may overlook some details so as to mistake like deviating from the dataset description. Inspired by recent studies about self-evaluation and self-alignment [54, 61, 62, 63, 16, 64], we leverage LLMs themselves to improve the quality of generated data. The process involves two primary steps: (1) Self-Reflection. Each generated data item is initially subjected to a self-reflection phase, wherein LLMs assess the item to determine errors and potential areas for enhancement. The output of self-reflection contains two parts: whether the given data needs to be enhanced and the reason why it needs enhancement. (2) Self-Enhancement. When LLMs recognize the necessity for improvements, both the reflective insights and the data item itself are re-input into the LLM to generate an improved version. By establishing a threshold for the number of iterations and repetitively applying these steps, UniGen effectively elevates the overall quality of the generated items.
Code-Based Mathematical Evaluation. In generating mathematics-related datasets, such as GSM8K [19], it has been observed that a proportion of generated labels are factually incorrect. To address this issue, we employ a code-based mathematical evaluation method to verify the accuracy of generated labels. As highlighted in the recent study by [65, 66], the use of tools (e.g., a Python function) can substantially improve reasoning performance. Motivated by this finding, we require the LLM to generate Python code to solve the given math-related problem. The code is then executed within a simulative environment to produce a solution. The code-verified answer(i.e., label) is subsequently compared with the original LLM-generated answer. If they conflict, the original LLM-generated answer will be replaced with the code-verified answer.
Truthfulness Validation by RAG. Ensuring the truthfulness of generated golden answers is crucial when creating datasets that require factual knowledge. Prior studies have utilized Retrieval-Augmented Generation (RAG) to enhance the factuality and reduce the incidence of hallucinations in LLMs [67, 68, 69, 70]. To combat hallucinations within the generated data, we implement a RAG-based validation process in UniGen. Specifically, the LLM first identifies keywords from the generated text. Subsequently, UniGen retrieves relevant descriptions based on these keywords from the Wikipedia database, as demonstrated in prior research [71]. These descriptions are then used as prompts to guide the LLM in detecting and correcting any discrepancies or errors in the generated content.

### v5

# DataGen: Unified Synthetic Dataset via Large Language Models

@@ Abstract
Large Language Models (LLMs) such as GPT-4 and Llama3 have significantly impacted various fields by enabling high-quality synthetic data generation and reducing dependence on expensive human-generated datasets. Despite this, challenges remain in the areas of generalization, controllability, diversity, and truthfulness within the existing generative frameworks. To address these challenges, this paper presents DataGen, a comprehensive LLM-powered framework designed to produce diverse, accurate, and highly controllable datasets. DataGen is adaptable, supporting all types of text datasets and enhancing the generative process through innovative mechanisms. To augment data diversity, DataGen incorporates an attribute-guided generation module and a group checking feature. For accuracy, it employs a code-based mathematical assessment for label verification alongside a retrieval-augmented generation technique for factual validation. The framework also allows for user-specified constraints, enabling customization of the data generation process to suit particular requirements. Extensive experiments demonstrate the superior quality of data generated by DataGen, and each module within DataGen plays a critical role in this enhancement. Additionally, DataGen is applied in two practical scenarios: benchmarking LLMs and data augmentation. The results indicate that DataGen effectively supports dynamic and evolving benchmarking and that data augmentation improves LLM capabilities in various domains, including agent-oriented abilities and reasoning skills. Code is available at https://github.com/HowieHwong/DataGen.
[…]
@@ 1 Introduction
Large Language Models (LLMs) such as GPT-4 (ChatGPT), Claude (Claude), and Llama3 (LLAMA3) have demonstrated excellent performance across various professional domains, including medical (liu2023deid; zhang2024biomedgpt), educational (kasneci2023chatgpt), software engineering (qian2023communicative), and social sciences (li2024i; li2024quantifying), as well as in LLM-based agent applications (huang2024metatool; liu2023agentbench; chen2024gui). Given their superior generative capabilities, it is natural for researchers to explore effective methods for utilizing these models in synthetic data generation (zhu2024dyval; zhu2024dyval2; wang2024benchmark). The primary goal is to produce high-quality, cost-effective datasets, thereby reducing the reliance on expensive human labor. Furthermore, data generated by LLMs can be utilized for data augmentation (yu2024large), dynamic evaluation (zhu2024dyval; zhu2024dyval2), and model self-alignment (sun2023principledriven).
Despite the advancements in LLM-driven data generation (zhu2024dyval; zhu2024dyval2; wang2024benchmark; dekoninck2024understanding; dekoninck2024controlled), which have significantly improved the data generation pipeline and reduced the human cost, some challenges remain: (1) Generalization and Controllability: Most of existing frameworks directly modify data items in original datasets in specific ways based on fixed principles (zhu2024dyval2; wang2024benchmark) (e.g., add additional context or shuffle the order of the options), which may constrain the generalization of the generated data as they do not modify the nature of the data items like the scenarios within items. Moreover, many of them are also limited to particular dataset formats or types (yu2024large; zhu2024dyval), such as multiple-choice or mathematically-oriented datasets (e.g., GSM8K (cobbe2021gsm8k)). Additionally, the lack of provisions for incorporating external constraints, like specific user requirements (e.g., users may specify the length of generated text), restricts their controllability during generation. (2) Diversity and Truthfulness: Prior efforts always overlook the need to ensure some quality aspects of the datasets like diversity and truthfulness. For instance, the direct application of LLMs for dataset generation often leads to replication and low diversity, as LLMs may output the same answers when faced with semantically similar input. Furthermore, the propensity of LLMs to produce hallucinations (huang2023survey; sun2024trustllm) can introduce factual inaccuracies, potentially degrading model performance when such datasets are used for training or fine-tuning.
Figure 1: Our proposed DataGen for dataset generation via LLMs.
@@
DyVal (zhu2024dyval) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
DyVal 2 (zhu2024dyval2) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
S3Eval (lei2024s3eval) $\checkmark$ $\checkmark$ $\checkmark$
yu2024large $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
chung2023increasing $\checkmark$ $\checkmark$
fan2024nphardeval $\checkmark$ $\checkmark$ $\checkmark$
jandaghi2023faithful $\checkmark$ $\checkmark$
wang2024benchmark $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MetaMath (yu2023metamath) $\checkmark$ $\checkmark$ $\checkmark$
Qameleon (agrawal2023qameleon) $\checkmark$
viswanathan2023prompt2model $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
chen-etal-2024-shot $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
gandhi-etal-2024-better $\checkmark$ $\checkmark$
DataGen (Ours) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
To address these challenges, this paper puts forward DataGen (as shown in Figure 1), a unified and LLM-powered framework designed to generate a dataset. DataGen ensures the generalization, diversity, truthfulness, and controllability simultaneously of the generation process, compared to previous studies (as shown in Table 1). DataGen accepts all kinds of text datasets and generates high-quality datasets based on various modules. To enrich the diversity of the generated datasets, DataGen employs a range of strategies, including various hyperparameter settings, attribute-guided generation, and group checking. To guarantee the truthfulness of the generated datasets, we propose a code-based mathematical assessment to detect and rectify potentially incorrect labels. Additionally, we adopt a Retrieval-Augmented Generation (RAG)-based validation method to check the factuality of generated statements to ensure their truthfulness. DataGen integrates constraints input to align with user specifications to enhance user control over the dataset generation process. Furthermore, by employing attribute-guided generation and difficulty enhancement, we enable the generation of data covering a wide range of topics while providing users with controllable difficulty levels.
@@
  • We introduce DataGen, a unified framework for generating textual datasets via LLMs, which accepts the original dataset, description, and user constraints, and integrates modules to ensure diversity, truthfulness, and controllability.
@@
  • We extensively evaluate DataGen across data characterization, module efficacy, human evaluation, error analysis, and cost analysis, confirming its proficiency in dataset generation and highlighting promising future research directions.
@@
  • We explore two applications of DataGen: benchmarking LLMs and data augmentation. Key insights include: I) Most LLMs struggle with math-oriented datasets generated by DataGen (e.g., GSM8K). II) The performance of LLMs varies significantly across datasets generated by different LLMs. III) LLMs’ capabilities across various aspects (e.g., agent-related abilities, reasoning skills) can be improved by fine-tuning based on the generated data. IV) An improvement of data augmentation exists in knowledge-intensive datasets.
2 DataGen Framework
Figure 2: The architecture of DataGen.
@@ 2 DataGen Framework
In this section, we will introduce the proposed DataGen, a unified framework for dataset generation. DataGen consists of four modules (as shown in Figure 2) including framework input, generation hint, internal evaluation, and post-processing. Formally, consider an original dataset $\mathcal{D}$, the proposed framework $\mathcal{F}$, which operates by iteratively sampling subsets $\mathcal{S}_{i}$ from $\mathcal{D}$ (i.e., example selection for few-shot learning in Section 2.2). For each subset, $\mathcal{F}$ applies transformations based on the dataset’s description $\mathcal{M}(\mathcal{D})$ and a set of constraints $\mathcal{C}$. The final generated dataset, $\mathcal{D}_{\text{gen}}$, is accumulated over $N$ iterations: $\mathcal{D}_{\text{gen}}=\bigcup_{i=1}^{N}\mathcal{F}(\mathcal{S}_{i},\mathcal{M}(\mathcal{D}),\mathcal{C})$. During generation, the objectives of DataGen focus on maximizing the generalization, controllability, diversity, and truthfulness of the generated dataset.
@@ 2.1 Framework Input
The input for DataGen comprises three components: base dataset, dataset description, and generation constraints: The base dataset is provided in a standardized JSON format, which may include text with a label or standalone text (e.g., “text with a label” or “single text”). The dataset description articulates the specifics of the base dataset at a high level, furnishing foundational guidance for the LLM to synthesize a dataset analogous to the original. While optional, the generation constraints (zhou2023instructionfollowing) specify fine-grained conditions under which the LLM operates during dataset generation. For instance, constraints might stipulate that “Do not generate text longer than 20 words” or “Include an emoji in each generated sample”, thereby restricting specific conditions of the synthetic dataset.
@@ 2.2 Generation Hint
Few-Shot Learning. The base dataset typically comprises hundreds of data items; however, incorporating all these items directly into the prompt may result in an excessively long context that could obscure the comprehension capabilities of LLMs and incur substantial costs (bai2023longbench). To mitigate these challenges, few-shot learning techniques are employed for dataset generation (brown2020language; wang2020generalizing). Within DataGen, two principal methods are utilized to select few-shot learning examples. The first method involves a random sampling from the base dataset, effectively reducing both generation time and associated costs. The second method focuses on enhancing the diversity of examples, thereby guiding LLMs to generate as varied a dataset as possible. Specifically, DataGen initially encodes all data items using OpenAI’s text-embedding-ada-002 (text-embedding-ada-002) to create an embedding list. Subsequently, a clustering algorithm (e.g., K-means (hartigan1979algorithm)) is applied to form $n$ clusters, where $n$ represents the desired number of examples. One example is randomly selected from each cluster, yielding a set of $n$ diverse examples.
Diversity Setting. To augment the diversity of the generated data, we implement two strategies: (1) Hyperparameter Setting. The content generated by LLMs is influenced by various factors, with hyperparameters such as temperature, top-k, and top-p being crucial. To maximize the diversity of the dataset, we manipulate these hyperparameters, particularly the temperature settings. (2) Attribute-Guided Generation. Drawing on insights from prior research (huang2024metatool; yu2024large), we formalize the attribute-guided text generation process for LLMs. Let $\mathcal{A}=\{a_{1},a_{2},\dots,a_{n}\}$ be a set of attributes, such as "economics" and "sports", intended to guide the generation process. We model the generation process as a function where the output text $y$ is a function of the input prompt $x$ and a vector of attributes $\mathbf{a}\in\mathcal{A}$. The generation process can be expressed as $y=P(x,\mathbf{a})$, where $P$ represents the generation model of the LLM, and $x$ is the input prompt. To implement this, we employ two distinct strategies: the first involves directly incorporating user-input customized attributes, and the second requires asking LLMs to extract necessary attributes from given data examples (the prompt template is shown in Appendix I). (3) Group Checking. To ensure diversity among the generated items, a similarity matrix is employed to identify and filter out pairs of data items exhibiting high similarity. Further details on this process are provided in Section 2.4.
@@ 2.3 Internal Evaluation
Overall Quality Assessment and Enhancement. After obtaining the raw generated data, it’s important to enhance their overall quality as during the generation, LLMs may overlook some details so as to mistake like deviating from the dataset description. Inspired by recent studies about self-evaluation and self-alignment (ji2023towards; ren2023selfevaluation; huang2023large; jain2023bring; sun2023principledriven; wang-etal-2023-self-instruct), we leverage LLMs themselves to improve the quality of generated data. The process involves two primary steps: (1) Self-Reflection. Each generated data item is initially subjected to a self-reflection phase, wherein LLMs assess the item to determine errors and potential areas for enhancement. The output of self-reflection contains two parts: whether the given data needs to be enhanced and the reason why it needs enhancement. (2) Self-Enhancement. When LLMs recognize the necessity for improvements, both the reflective insights and the data item itself are re-input into the LLM to generate an improved version. By establishing a threshold for the number of iterations and repetitively applying these steps, DataGen effectively elevates the overall quality of the generated items.
Code-Based Mathematical Evaluation. In generating mathematics-related datasets, such as GSM8K (cobbe2021gsm8k), it has been observed that a proportion of generated labels are factually incorrect. To address this issue, we employ a code-based mathematical evaluation method to verify the accuracy of generated labels. As highlighted in the recent study by (gou2024tora; chen2023fireact), the use of tools (e.g., a Python function) can substantially improve reasoning performance. Motivated by this finding, we require the LLM to generate Python code to solve the given math-related problem. The code is then executed within a simulative environment to produce a solution. The code-verified answer(i.e., label) is subsequently compared with the original LLM-generated answer. If they conflict, the original LLM-generated answer will be replaced with the code-verified answer.
Truthfulness Validation by RAG. Ensuring the truthfulness of generated golden answers is crucial when creating datasets that require factual knowledge. Prior studies have utilized Retrieval-Augmented Generation (RAG) to enhance the factuality and reduce the incidence of hallucinations in LLMs (aksitov2023rest; li2024enhancing; li2022decoupled; gao2024best). To combat hallucinations within the generated data, we implement a RAG-based validation process in DataGen. Specifically, the LLM first identifies keywords from the generated text. Subsequently, DataGen retrieves relevant descriptions based on these keywords from the Wikipedia database, as demonstrated in prior research (semnani2023wikichat). These descriptions are then used as prompts to guide the LLM in detecting and correcting any discrepancies or errors in the generated content.